# Model Reveals What to Cache: Profiling-Based Feature Reuse for Video Diffusion Models

### v1

@@ 1 Introduction
Video generation has emerged as a pivotal technology in computer vision, driving advancements across diverse domains from multimedia applications and entertainment to interactive systems [27, 15, 6, 36, 25, 21, 2]. This surge in demand has driven rapid progress in generative modeling [33, 3, 29], particularly diffusion-based approaches. Among these methods, the Diffusion Transformer (DiT) architecture [26] has gained prominence for its ability to synthesize high-fidelity and temporally consistent videos, supporting downstream tasks including augmented reality [8], automated content creation [37, 34], and AI-driven storytelling [41, 10, 39]. However, despite their remarkable success in visual quality and diversity, DiT models face a critical challenge: their inherently sequential denoising process necessitates iterative computations, resulting in substantial computational overhead and substantial memory demands, particularly for high-resolution video generation and large-scale models.
@@ 3.1 Preliminaries
Figure 3: Heatmap of attention scores. Attention scores are displayed at the block level (rows) and step level (columns), revealing consistent patterns across steps. Red denotes regions with the highest attention, followed by green and blue, while purple indicates areas with the lowest attention.
@@ 3.2 Attention Pattern Analysis
Given an input prompt, such as ”a cat walking on grass”, we analyze the attention behavior within DiT blocks. Each block contains a self-attention layer, from which token embeddings $\mathbf{X}\in\mathbb{R}^{B\times N\times C}$ are extracted across all denoising steps. The token dimension $N$ is divided into $T$ segments, each corresponding to one video frame. Attention scores are computed for each frame by summing across rows in the attention matrix.
Block-Level Analysis. To determine whether each block predominantly attends to foreground or background regions, we first extract the noise prediction at each step (shape: $B\times C\times T\times H\times W$). To segment the latent space into foreground and background, we apply PCA [1] to reduce the channel dimension to 3, forming RGB-like visualizations as shown in Fig. 2. We then use SAM2 to generate binary masks of the foreground.
@@
where $N_{\text{high}\cap\text{fg}}$ denotes the number of tokens that overlap between high-attention$(\bar{a}_{i}>\text{threshold})$ tokens and the foreground mask, and $N_{\text{fg}}$ is the total number of foreground tokens. Then we can get a block list from: We define the foreground ($\mathcal{F}$) and background ($\mathcal{B}$) blocks list based on a predefined threshold $\tau$:
@@ 3.3 ProfilingDiT
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (ms) $\downarrow$ Speedup $\uparrow$
@@
Table 1: Quantitative comparison with prior methods under HunyuanVideo baselines. “Ours” consistently outperforms others in visual quality while maintaining strong efficiency. $\uparrow$: higher is better; $\downarrow$: lower is better.
@@
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (ms) $\downarrow$ Speedup $\uparrow$
@@
Table 2: Quantitative comparison with prior methods under Wan2.1 baselines.
@@
However, based on our previous observations, the block focus on the background list contains both consecutive numbers and individual numbers. The Delta-DiT only caches a single group of consecutive blocks, but we need to process all the blocks in the list. Therefore, we have designed the following delta list scheme, as shown in Fig. 6.
@@
where $f(s)$ is a monotonically decreasing function, ensuring that caching occurs less frequently at the beginning and more frequently in later steps.
@@ 4 Experiment
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (ms) $\downarrow$ Speedup $\uparrow$
@@
Table 3: Ablation on block-level reuse strategies. Foreground reuse achieves the best perceptual quality and distortion metrics. “Split” alternates block types across step groups, while “Alternate” switches reuse types between steps.
@@
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (ms) $\downarrow$ Speedup $\uparrow$
@@
Table 4: Ablation on step-level reuse strategies across background and foreground blocks. ”Stepwise” denotes decreasing reuse intervals; ”Step Inverse” denotes increasing reuse intervals; ”Step Average” denotes fixed reuse intervals.
@@ 4.4 Ablation Study
We first investigate the impact of reusing different types of blocks—Foreground-focused versus Background-focused. The results are shown in Tab. 3. Our analysis shows that Background-focused block reuse (i.e., reusing blocks that focus on background content) helps preserve global consistency and yields a competitive VBench score (0.768), but struggles with fine-grained visual quality (LPIPS = 0.1449, PSNR = 24.23, FID = 58.53). In contrast, Foreground-focused block reuse (i.e., reusing blocks that focus on dynamic foreground content) is more effective for detailed reconstruction, achieving substantially better LPIPS (0.1203), PSNR (26.44), and SSIM (0.8445), along with a lower FID of 41.10. These results suggest that background features are more spatially localized and temporally variant, thus benefiting more from selective reuse.
@@
To explore how reuse frequency affects performance, we vary the reuse interval, defined as the number of steps a cached block is retained before recomputation. As shown in Tab. 4, a “Stepwise” strategy—where reuse intervals gradually decrease—consistently improves perceptual quality over fixed (“Step Average”) or increasing (“Step Inverse”) schedules. For instance, the Background+Stepwise strategy outperforms its inverse counterpart in both PSNR (26.44 vs. 24.49) and LPIPS (0.1203 vs. 0.2529), while being nearly twice as fast as the baseline (932 ms vs. 1745 ms). These findings highlight that reusing stable (background-focused) blocks early, followed by more frequent recomputation of dynamic (foreground-focused) blocks later, helps avoid stale information accumulation while maintaining quality.

### v2

@@ 1 Introduction
Video generation has emerged as a pivotal technology in computer vision, driving advancements across diverse domains from multimedia applications and entertainment to interactive systems [27, 15, 6, 38, 25, 21, 2]. This surge in demand has driven rapid progress in generative modeling [35, 3, 31], particularly diffusion-based approaches. Among these methods, the Diffusion Transformer (DiT) architecture [26] has gained prominence for its ability to synthesize high-fidelity and temporally consistent videos, supporting downstream tasks including augmented reality [8], automated content creation [39, 36, 30], and AI-driven storytelling [43, 10, 41]. However, despite their remarkable success in visual quality and diversity, DiT models face a critical challenge: their inherently sequential denoising process necessitates iterative computations, resulting in substantial computational overhead and substantial memory demands, particularly for high-resolution video generation and large-scale models.
@@ 3.1 Preliminaries
Figure 3: Visualization of the attention scores heatmap. Attention scores are displayed at the block level (rows) and step level (columns), revealing consistent patterns across steps. Red denotes regions with the highest attention, followed by green and blue, while purple indicates areas with the lowest attention.
@@ 3.2 Attention Pattern Analysis
Given an input prompt, such as “a cat walking on grass”, we analyze the attention behavior within DiT blocks. Each block contains a self-attention layer, from which token embeddings $\mathbf{X}\in\mathbb{R}^{B\times N\times C}$ are extracted across all denoising steps. The token dimension $N$ is divided into $T$ segments, each corresponding to one video frame. Attention scores are computed for each frame by summing across rows in the attention matrix.
Block-Level Analysis. To determine whether each block predominantly attends to foreground or background regions, we first extract the noise prediction at each step (shape: $B\times C\times T\times H\times W$). To segment the latent space into foreground and background, we apply PCA [1] to reduce the channel dimension to 3, forming RGB-like visualizations as shown in Fig. 2. We then use SAM2 [28] to generate binary masks of the foreground.
@@
[…]
@@ 3.3 ProfilingDiT
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (s) $\downarrow$ Speedup $\uparrow$
@@
Table 1: Quantitative comparison with prior methods under HunyuanVideo baselines. “Ours” consistently outperforms others in visual quality while maintaining strong efficiency. $\uparrow$: higher is better; $\downarrow$: lower is better. Bolding and underlining indicate the best and second-best performance, respectively.
@@
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (s) $\downarrow$ Speedup $\uparrow$
@@
Table 2: Quantitative comparison with prior methods under Wan2.1 baselines. Bolding and underlining indicate the best and second-best performance, respectively.
@@
However, based on our previous observations, the block focuses on the background list that contains both consecutive numbers and individual numbers. The Delta-DiT only caches a single group of consecutive blocks, but we need to process all the blocks in the list. Therefore, we have designed the following delta list scheme, as shown in Fig. 6.
@@
where $f(s)$ is a monotonically decreasing function representing the caching interval. For example, the intervals can be $f(s)\in\{12,9,6,3\}$, indicating that caching becomes progressively more frequent over time.
@@ 4 Experiment
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (s) $\downarrow$ Speedup $\uparrow$
@@
Table 3: Ablation on block-level reuse strategies. Foreground reuse achieves the best perceptual quality and distortion metrics. “Split” alternates block types across step groups, while “Alternate” switches reuse types between steps. Bolding and underlining indicate the best and second-best performance, respectively.
@@
VBench $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ FID $\downarrow$ Latency (s) $\downarrow$ Speedup $\uparrow$
@@
Table 4: Ablation on step-level reuse strategies across background and foreground blocks. “Stepwise” denotes decreasing reuse intervals; “Step Inverse” denotes increasing reuse intervals; “Step Average” denotes fixed reuse intervals. Bolding and underlining indicate the best and second-best performance, respectively.
@@ 4.4 Ablation Study
We first investigate the impact of reusing different types of blocks—Foreground-focused versus Background-focused. The results are shown in Tab. 3. Our analysis shows that Foreground-focused block reuse (i.e., reusing blocks that focus on foreground content) yields a competitive VBench score (0.768), but struggles with fine-grained visual quality (LPIPS = 0.1449, PSNR = 24.23, FID = 58.53). In contrast, Background-focused block reuse (i.e., reusing blocks that focus on dynamic background content) is more effective for detailed reconstruction, achieving substantially better LPIPS (0.1203), PSNR (26.44), and SSIM (0.8445), along with a lower FID of 41.10. These results suggest that background features are more spatially localized and temporally variant, thus benefiting more from selective reuse.
@@
To explore how reuse frequency affects performance, we vary the reuse interval, defined as the number of steps a cached block is retained before recomputation. As shown in Tab. 4, a “Stepwise” strategy—where reuse intervals gradually decrease—consistently improves perceptual quality over fixed (“Step Average”) or increasing (“Step Inverse”) schedules. For instance, the Background+Stepwise strategy outperforms its inverse counterpart in both PSNR (26.44 vs. 24.49) and LPIPS (0.1203 vs. 0.2529), while being nearly twice as fast as the baseline (932 s vs. 1745 s). These findings highlight that reusing stable (background-focused) blocks early, followed by more frequent recomputation of dynamic (foreground-focused) blocks later, helps avoid stale information accumulation while maintaining quality.